# Beyond algorithm hyperparameters: on preprocessing hyperparameters and associated pitfalls in machine learning applications

### v1

@@ 3.2.1 Evaluation of a model generated on all available data
A straightforward way to evaluate a prediction model trained on all available data is to estimate its prediction error using the same data set, i.e. $\mathcal{D}_{\text{train}}=\mathcal{D}_{\text{test}}=\mathcal{D}$. The resulting prediction error estimate is referred to as apparent error (see Figure 2, model evaluation a). As explained in Section 2.4.2, data leakage is present when information about the designated $\mathcal{D}_{\text{test}}$ is present during model generation. For the apparent error, this is clearly the case, as $\mathcal{D}_{\text{test}}$ is equal to $\mathcal{D}_{\text{train}}$. As a consequence, the apparent error is not able to detect any overfitting of the model (since the specific pattern of noise in $\mathcal{D}_{\text{train}}$ exactly corresponds to that in $\mathcal{D}_{\text{test}}$) and will therefore be affected by a (possibly substantial) optimistic bias. Although this evaluation procedure is well-known to be flawed and has been frequently warned against in literature (e.g., Efron,, 1986; Kuhn and Johnson,, 2013; Hastie et al.,, 2009; Poldrack et al.,, 2020; Collins et al., 2024a, ), it is often still the only prediction error estimate that is reported in studies presenting new prediction models (Kapoor and Narayanan,, 2023; Poldrack et al.,, 2020).
@@
To avoid the optimistic bias caused by the overlap between $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$, several procedures exist that partition $\mathcal{D}_{\text{train}}$ one or multiple times into two subsets for evaluation purposes, while still training the final prediction model on the full data set. These procedures can be referred to as resampling methods and the resulting estimate as resampling error (see Figure 2, model evaluation b). The following description is based on Simon, (2007), Kuhn and Johnson, (2013), Bischl et al., (2023), and Casalicchio and Burk, (2024); see their work for more details.
@@
A commonly-used variation of holdout is $k$-fold cross-validation (CV), where $\mathcal{D}_{\text{train}}$ is randomly split into $k$ subsets (or folds) of approximately the same size, with 5 or 10 being typical choices for $k$. Based on the $k$ splits, the procedure described for the holdout method is repeated $k$ times: In each repetition (in this context also referred to as resampling iteration), the learning pipeline is trained on $k-1$ subsets of $\mathcal{D}_{\text{train}}$ (constituting $\mathcal{D}^{\prime}_{\text{train}}$), and the prediction error of the resulting model is estimated on the remaining subset (constituting $\mathcal{D}_{\text{test}}$). The final prediction error estimate is obtained by averaging the $k$ prediction error estimates, which leads to the CV estimator having a smaller variance than a holdout estimator with the same split ratio. However, the prediction error estimate resulting from CV is also pessimistically biased because the evaluated prediction models are again trained on less than $n_{\text{train}}$ observations, although this bias decreases with increasing $k$ ($n^{\prime}_{\text{train}}=\frac{k-1}{k}\cdot n_{\text{train}}$).
@@
As stated above, all resampling methods require the learning pipeline to be retrained on one or multiple subsets $\mathcal{D}^{\prime}_{\text{train}}$, each of which is a (proper) subset of $\mathcal{D}_{\text{train}}$ (i.e. $\mathcal{D}^{\prime}_{\text{train}}\subset\mathcal{D}_{\text{train}}$). In this context, a flawed evaluation procedure would be to apply all preprocessing steps on the full data set $\mathcal{D}_{\text{train}}$ and retrain only the learning algorithm on $\mathcal{D}^{\prime}_{\text{train}}$ during resampling. This “incomplete resampling” (Simon et al.,, 2003) results in another form of data leakage, as in each resampling iteration, the observations in the respective $\mathcal{D}_{\text{test}}$ subset have already been used to train part of the learning pipeline (i.e. the preprocessing steps). Incomplete resampling has been frequently warned against in the literature (e.g., Kapoor et al.,, 2024; Poldrack et al.,, 2020; Pfob et al.,, 2022; de Hond et al.,, 2022; Hofman et al.,, 2023), and the resulting optimistic bias has been demonstrated by illustrations on real data (e.g., Hornung et al.,, 2015; Rosenblatt et al.,, 2024) and corrected reanalyses of published studies (e.g., Neunhoeffer and Sternberg,, 2019; Kapoor and Narayanan,, 2023). Yet, it still seems to be a common pitfall in the evaluation of prediction models (see Kapoor and Narayanan,, 2023 and references therein), which is probably caused by a lack of understanding of its implications. In addition, if the learning pipeline is not implemented as a single object that can be trained with a single function call such as train(learning_pipeline) (e.g., this is possible in R with the mlr3 or recipes package by Lang et al.,, 2019 and Kuhn et al.,, 2024), each preprocessing step must be manually repeated in every resampling iteration. In such cases, users may consider incomplete resampling a time-saving shortcut, without realizing that it introduces data leakage. To avoid incomplete resampling, every component of the learning pipeline, including the preprocessing steps, must be retrained in each resampling iteration. The only preprocessing steps that can be safely applied to the full data set prior to resampling are those that are both parameterless and precede the first parameterized preprocessing step in the learning pipeline.
@@ 3.2.2 Evaluation of a model generated on a subset of the available data
In principle, most points discussed in the previous section affecting temporary holdout (including data leakage due to incomplete resampling) also apply to permanent holdout. Again, the only difference is that, for the temporary holdout, the model trained on a subset of the available data is used solely for evaluation purposes, whereas it serves as the final prediction model for the permanent holdout. Consequently, the prediction error estimate derived from the permanent holdout is not pessimistically biased; instead, it is an unbiased estimate of a prediction error that is indeed higher (i.e. worse) than that of a model using all available data. Since not using all available data for training the prediction model essentially corresponds to a loss of important information, the permanent holdout method is only recommended if the number of observations in $\mathcal{D}$ is sufficiently large, or if repeating the training process is computationally expensive or infeasible (Collins et al., 2024a, ).
@@ 4.1.1 Overview
HP tuning generally aims to improve the predictive performance of a model (Bischl et al.,, 2023; Probst et al.,, 2019). Using the terminology introduced in Section 2.4.2, this corresponds to finding the HP configuration that minimizes the model’s prediction error. To simplify notation, we will assume for now that all HPs are to be tuned, but will revisit the scenario where this does not apply later in this section. Under this assumption, the HP tuning problem can be formalized as:
@@
where $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}_{\boldsymbol{\lambda}}}$ is the final prediction model resulting from training the learning pipeline $\mathcal{I}$ configured with HPs $\boldsymbol{\lambda}$, and $\boldsymbol{\lambda}^{\ast}$ denotes the theoretical optimum (Bischl et al.,, 2023). The lowest prediction error (i.e. the best performance) that can be achieved using $\boldsymbol{\lambda}^{\ast}$ as HP configuration depends on several factors, such as the HPs to be tuned, the selected learning algorithm, the performance measure, and the prediction problem in general (Probst et al.,, 2019). Note that in the following, we refer to the prediction error of a model that results from training a learning pipeline determined by a candidate HP configuration $\boldsymbol{\lambda}^{(c)}$, i.e. $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}_{\boldsymbol{\lambda}^{(c)}}}$, simply as the prediction error of $\boldsymbol{\lambda}^{(c)}$ for brevity. It should also be noted that equation (3) represents the standard case of single-objective HP tuning, i.e. the optimization is performed with respect to one performance measure. However, HP tuning can also be conducted based on multiple performance measures or additional criteria such as model simplicity (Bischl et al.,, 2023; Dunias et al.,, 2024). Since such multi-objective HP tuning poses further challenges, we will only consider single-objective tuning in this paper.
@@
While there exist different tuning procedures, the general model generation process involving tuning can be described as follows: given a set of $C$ candidate HP configurations (selected before or during the tuning process), each HP configuration $\boldsymbol{\lambda}^{(c)}$ ($c=1,\dots,C$) is evaluated on $\mathcal{D}_{\text{train}}$ by employing one of the model evaluation procedures introduced in Section 3.2.1. Accordingly, $\mathcal{D}_{\text{train}}$ is split into $\mathcal{D}^{\prime}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ (either once or multiple times), which are then used for training ($\mathcal{D}^{\prime}_{\text{train}}$) and prediction error estimation ($\mathcal{D}_{\text{test}}$). In other words, the model evaluation that is performed once with $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{\mathrm{I}}$ in Setting I to assess the prediction error of the final prediction model is performed multiple times for each candidate configuration (i.e. with $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{(c)}$) in the tuning process of Setting II. After having evaluated all candidate HP configurations, the HP configuration with the lowest (i.e. best) prediction error estimate is used as the final HP configuration. Following the notation introduced in Section 2.3.4, we refer to this configuration as $\boldsymbol{\lambda}^{\mathrm{II}}$. Note that $\boldsymbol{\lambda}^{\mathrm{II}}$ is also commonly denoted as $\hat{\boldsymbol{\lambda}}$, since it is an estimate of $\boldsymbol{\lambda}^{\ast}$ (Bischl et al.,, 2023). However, we adhere to $\boldsymbol{\lambda}^{\mathrm{II}}$ to clearly distinguish it from Setting I, where $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{\mathrm{I}}$. After setting $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{\mathrm{II}}$, the learning pipeline $\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{II}}}$ undergoes a final training on $\mathcal{D}_{\text{train}}$, which yields the final prediction model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{II}}}}$.
@@
Note that while the tuning process already results in a prediction error estimate for the final prediction model (the estimate based on which $\boldsymbol{\lambda}^{\mathrm{II}}$ was selected during tuning) this value is not necessarily adopted as the final model evaluation result, as we will discuss in Section 4.2. In fact, it is also possible to use different performance measures for the prediction error estimation performed during tuning and the evaluation of the final model, but, for the sake of simplicity, we will assume that they are the same.
@@
To summarize, during the model generation in Setting II, both the HPs $\boldsymbol{\lambda}$ and the parameters $\boldsymbol{\theta}$ of the final prediction model are optimized using the data set $\mathcal{D}_{\text{train}}$. However, the optimization is not performed jointly: first, the HPs $\boldsymbol{\lambda}$ are optimized in the tuning process. Second, the parameters $\boldsymbol{\theta}$ are optimized in one (final) training process. Note that HPs are still an input of the learning pipeline, but can be seen as an output of the tuning process.
@@
If only a subset of the HPs $\boldsymbol{\lambda}$ are to be tuned, the tuning process described above is applied exclusively to those HPs, while the pre-specified HPs remain fixed throughout the process. For example, assume that from all $J$ HPs in $\boldsymbol{\lambda}$, the HPs $\boldsymbol{\lambda}_{1:j}=\lambda_{1},..\lambda_{j}$ are pre-specified and the HPs $\boldsymbol{\lambda}_{j+1:J}=\lambda_{j+1},\dots,\lambda_{J}$ are to be tuned. In this case, the tuning process yields a HP configuration $\boldsymbol{\lambda}_{j+1:J}^{\mathrm{II}}$, and the final prediction model is trained with $\boldsymbol{\lambda}_{1:j}=\boldsymbol{\lambda}_{1:j}^{\mathrm{I}}$ and $\boldsymbol{\lambda}_{j+1:J}=\boldsymbol{\lambda}_{j+1:J}^{\mathrm{II}}$. Since the tuning process is conceptually the same when not all HPs are optimized – untuned HPs are simply kept fixed – we will continue to assume that all HPs are tuned to maintain notational simplicity.
@@
When choosing a tuning procedure, it is important to consider that the tuning process is limited in terms of both data availability and computation time: First, as outlined above, each candidate HP configuration, $\boldsymbol{\lambda}^{(c)}$, is evaluated using one of the evaluation procedures described in Section 3.2.1 for Setting I. As explained there, the specified $\mathcal{D}^{\prime}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ subsets contain a limited number of observations (i.e. $n^{\prime}_{\text{train}}$ and $n_{\text{test}}$ $\leq n_{\text{train}}$) and could overlap, potentially leading to unreliable prediction error estimates for each $\boldsymbol{\lambda}^{(c)}$. Second, the computational budget available for the tuning process is typically limited, which restricts both the number of evaluated HP configurations and the time spent evaluating each configuration (i.e. estimating its prediction error). Due to these limitations and the resulting trade-offs (discussed in more detail in Section 4.1.3), choosing an adequate tuning procedure is often non-trivial. Yet, guidance is still lacking and many of the existing recommendations are based on rules of thumb rather than empirical benchmarks (see Bischl et al.,, 2023 for an overview). Inadequate tuning procedures can result in a $\boldsymbol{\lambda}^{\mathrm{II}}$ that yields a final prediction model with worse prediction error than $\boldsymbol{\lambda}^{\ast}$ (potentially even worse than setting all HPs to their default values) and/or an overly time-consuming tuning process (i.e. a more efficient tuning procedure could have achieved the same prediction error in less time).
@@
Before describing different tuning procedures in more detail, we note that their specification generally depends on whether the tuning process is fully automated or performed manually. We consider the tuning process as automated if the relevant tuning components only need to be specified as a function argument, which is possible in several ML software frameworks (see Bischl et al.,, 2023 for an overview). In contrast, we refer to the tuning process as manual if the candidate HP configurations are evaluated by repeatedly calling the same function(s), altering only the argument that specifies the HP configuration.
Compared to automated tuning, manual tuning is more time-consuming, error-prone, and less reproducible, as it is usually an informal and unsystematic process. On the other hand, automated tuning is usually more difficult to implement and requires more programming expertise than manual tuning. As a consequence, although manual tuning is generally advised against (e.g., Bartz et al.,, 2023; Bischl et al.,, 2023), it is likely still a common yet often unreported approach in many ML applications (Hosseini et al.,, 2020; Hofman et al.,, 2023; Lones,, 2024). Note that this may be particularly true for the tuning of preprocessing HPs $\boldsymbol{\lambda}_{P}$: As discussed in Section 2.3.3, preprocessing HPs are often not identified as HPs. Consequently, users trying out different preprocessing options might not be aware that this corresponds to (manual) HP tuning and could be automated. Moreover, if the HPs to be tuned include application-specific preprocessing HPs, the barrier to using automated tuning is further increased, as these HPs may not yet be integrated into the corresponding software and require custom implementation.

### v2

@@ 3.2.1 Evaluation of a model generated on all available data
A straightforward way to evaluate a prediction model trained on all available data is to estimate its prediction error using the same data set, i.e. $\mathcal{D}_{\text{train}}=\mathcal{D}_{\text{test}}=\mathcal{D}$. The resulting prediction error estimate is referred to as apparent error (see Figure 2, model evaluation a). As explained in Section 2.4.2, data leakage is present when information about the designated $\mathcal{D}_{\text{test}}$ is present during model generation. For the apparent error, this is clearly the case, as $\mathcal{D}_{\text{test}}$ is equal to $\mathcal{D}_{\text{train}}$. As a consequence, the apparent error is not able to detect any overfitting of the model (since the specific pattern of noise in $\mathcal{D}_{\text{train}}$ exactly corresponds to that in $\mathcal{D}_{\text{test}}$) and will therefore be affected by a (possibly substantial) optimistic bias. Although this evaluation procedure is well-known to be flawed and has been frequently warned against in literature [93, 111, 98, 122, 84, e.g.,], it is often still the only prediction error estimate that is reported in studies presenting new prediction models [108, 122].
@@
To avoid the optimistic bias caused by the overlap between $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$, several procedures exist that partition $\mathcal{D}_{\text{train}}$ one or multiple times into two subsets for evaluation purposes while still training the final prediction model on the full data set. These procedures can be referred to as resampling methods and the resulting estimate as the resampling error (see Figure 2, model evaluation b). The following description is based on [131], [111], [78], and [82]; see their work for more details.
@@
A commonly used variation of holdout is $k$-fold cross-validation (CV), where $\mathcal{D}_{\text{train}}$ is randomly split into $k$ subsets (or folds) of approximately the same size, with 5 or 10 being typical choices for $k$. Based on the $k$ splits, the procedure described for the holdout method is repeated $k$ times: In each repetition (in this context also referred to as resampling iteration), the learning pipeline is trained on k1k-1 subsets of $\mathcal{D}_{\text{train}}$ (constituting $\mathcal{D}^{\prime}_{\text{train}}$), and the prediction error of the resulting model is estimated on the remaining subset (constituting $\mathcal{D}_{\text{test}}$). The final prediction error estimate is obtained by averaging the $k$ prediction error estimates, which leads to the CV estimator having a smaller variance than a holdout estimator with the same split ratio. However, the prediction error estimate resulting from CV is also pessimistically biased because the evaluated prediction models are again trained on less than $n_{\text{train}}$ observations, although this bias decreases with increasing $k$ ($n^{\prime}_{\text{train}}=\frac{k-1}{k}\cdot n_{\text{train}}$).
@@
As stated above, all resampling methods require the learning pipeline to be retrained on one or multiple subsets $\mathcal{D}^{\prime}_{\text{train}}$, each of which is a (proper) subset of $\mathcal{D}_{\text{train}}$ (i.e. $\mathcal{D}^{\prime}_{\text{train}}\subset\mathcal{D}_{\text{train}}$). In this context, a flawed evaluation procedure would be to apply all preprocessing steps on the full data set $\mathcal{D}_{\text{train}}$ and retrain only the learning algorithm on $\mathcal{D}^{\prime}_{\text{train}}$ during resampling. This “incomplete resampling” [130] results in another form of data leakage, as in each resampling iteration, the observations in the respective $\mathcal{D}_{\text{test}}$ subset have already been used to train part of the learning pipeline (i.e. the preprocessing steps). Incomplete resampling has been frequently warned against in the literature [107, 122, 121, 86, 100, e.g.,], and the resulting optimistic bias has been demonstrated by illustrations on real data [102, 127, e.g.,] and corrected reanalyses of published studies [119, 108, e.g.,]. Yet, it still seems to be a common pitfall in the evaluation of prediction models (see [108], and references therein), which is probably caused by a lack of understanding of its implications. In addition, if the learning pipeline is not implemented as a single object that can be trained with a single function call such as train(learning_pipeline) (e.g., this is possible in R with the mlr3 or recipes package by [113], and [112]), each preprocessing step must be manually repeated in every resampling iteration. In such cases, users may consider incomplete resampling a time-saving shortcut, without realizing that it introduces data leakage. To avoid incomplete resampling, every component of the learning pipeline, including the preprocessing steps, must be retrained in each resampling iteration. The only preprocessing steps that can be safely applied to the full data set prior to resampling are those that are both parameterless and precede the first parameterized preprocessing step in the learning pipeline.
@@ 3.2.2 Evaluation of a model generated on a subset of the available data
In principle, most points discussed in the previous section affecting temporary holdout (including data leakage due to incomplete resampling) also apply to permanent holdout. Again, the only difference is that, for the temporary holdout, the model trained on a subset of the available data is used solely for evaluation purposes, whereas it serves as the final prediction model for the permanent holdout. Consequently, the prediction error estimate derived from the permanent holdout is not pessimistically biased; instead, it is an unbiased estimate of a prediction error that is indeed higher (i.e. worse) than that of a model using all available data. Since not using all available data for training the prediction model essentially corresponds to a loss of important information, the permanent holdout method is only recommended if the number of observations in $\mathcal{D}$ is sufficiently large or if repeating the training process is computationally expensive or infeasible [84].
@@ 4.1.1 Overview
HP tuning generally aims to improve the predictive performance of a model [78, 124]. Using the terminology introduced in Section 2.4.2, this corresponds to finding the HP configuration that minimizes the model’s prediction error. To simplify notation, we will assume for now that all HPs are to be tuned, but will revisit the scenario where this does not apply later in this section. Under this assumption, the HP tuning problem can be formalized as:
@@
where $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}_{\boldsymbol{\lambda}}}$ is the final prediction model resulting from training the learning pipeline $\mathcal{I}$ configured with HPs $\boldsymbol{\lambda}$, and $\boldsymbol{\lambda}^{\ast}$ denotes the theoretical optimum [78]. The lowest prediction error (i.e. the best performance) that can be achieved using $\boldsymbol{\lambda}^{\ast}$ as HP configuration depends on several factors, such as the HPs to be tuned, the selected learning algorithm, the performance measure, and the prediction problem in general [124]. Note that in the following, we refer to the prediction error of a model that results from training a learning pipeline determined by a candidate HP configuration $\boldsymbol{\lambda}^{(c)}$, i.e. $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}_{\boldsymbol{\lambda}^{(c)}}}$, simply as the prediction error of $\boldsymbol{\lambda}^{(c)}$ for brevity. It should also be noted that equation (3) represents the standard case of single-objective HP tuning, i.e. the optimization is performed with respect to one performance measure. However, HP tuning can also be conducted based on multiple performance measures or additional criteria such as model simplicity [78, 91]. Since such multi-objective HP tuning poses further challenges, we will only consider single-objective tuning in this paper.
@@
While there exist different tuning procedures, the general model generation process involving tuning can be described as follows: Given a set of $C$ candidate HP configurations (selected before or during the tuning process), each HP configuration $\boldsymbol{\lambda}^{(c)}$ ($c=1,\dots,C$) is evaluated on $\mathcal{D}_{\text{train}}$ by employing one of the model evaluation procedures introduced in Section 3.2.1. Accordingly, $\mathcal{D}_{\text{train}}$ is split into $\mathcal{D}^{\prime}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ (either once or multiple times), which are then used for training ($\mathcal{D}^{\prime}_{\text{train}}$) and prediction error estimation ($\mathcal{D}_{\text{test}}$). In other words, the model evaluation that is performed once with $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{\mathrm{I}}$ in Setting I to assess the prediction error of the final prediction model is performed multiple times for each candidate configuration (i.e. with $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{(c)}$) in the tuning process of Setting II. After having evaluated all candidate HP configurations, the HP configuration with the lowest (i.e. best) prediction error estimate is used as the final HP configuration. Following the notation introduced in Section 2.3.4, we refer to this configuration as $\boldsymbol{\lambda}^{\mathrm{II}}$. Note that $\boldsymbol{\lambda}^{\mathrm{II}}$ is also commonly denoted as $\hat{\boldsymbol{\lambda}}$, since it is an estimate of $\boldsymbol{\lambda}^{\ast}$ [78]. However, we adhere to $\boldsymbol{\lambda}^{\mathrm{II}}$ to clearly distinguish it from Setting I, where $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{\mathrm{I}}$. After setting $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{\mathrm{II}}$, the learning pipeline $\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{II}}}$ undergoes a final training on $\mathcal{D}_{\text{train}}$, which yields the final prediction model $\hat{f}^{\mathcal{D}_{\text{train}}}_{\mathcal{I}_{\boldsymbol{\lambda}^{\mathrm{II}}}}$.
@@
Note that while the tuning process already results in a prediction error estimate for the final prediction model (the estimate based on which $\boldsymbol{\lambda}^{\mathrm{II}}$ was selected during tuning), this value is not necessarily adopted as the final model evaluation result, as we will discuss in Section 4.2. In fact, it is also possible to use different performance measures for the prediction error estimation performed during tuning and the evaluation of the final model, but, for the sake of simplicity, we will assume that they are the same.
@@
To summarize, during the model generation in Setting II, both the HPs $\boldsymbol{\lambda}$ and the parameters $\boldsymbol{\theta}$ of the final prediction model are optimized using the data set $\mathcal{D}_{\text{train}}$. However, the optimization is not performed jointly: first, the HPs $\boldsymbol{\lambda}$ are optimized in the tuning process. Second, the parameters $\boldsymbol{\theta}$ are optimized in one (final) training process. Note that HPs are still an input of the learning pipeline but can be seen as an output of the tuning process.
@@
If only a subset of the HPs $\boldsymbol{\lambda}$ are to be tuned, the tuning process described above is applied exclusively to those HPs, while the pre-specified HPs remain fixed throughout the process. For example, assume that from all $J$ HPs in $\boldsymbol{\lambda}$, the HPs $\boldsymbol{\lambda}_{1:j}=\lambda_{1},..\lambda_{j}$ are pre-specified and the HPs $\boldsymbol{\lambda}_{j+1:J}=\lambda_{j+1},\dots,\lambda_{J}$ are to be tuned. In this case, the tuning process yields a HP configuration $\boldsymbol{\lambda}_{j+1:J}^{\mathrm{II}}$, and the final prediction model is trained with $\boldsymbol{\lambda}_{1:j}=\boldsymbol{\lambda}_{1:j}^{\mathrm{I}}$ and $\boldsymbol{\lambda}_{j+1:J}=\boldsymbol{\lambda}_{j+1:J}^{\mathrm{II}}$. Since the tuning process is conceptually the same when not all HPs are optimized—untuned HPs are simply kept fixed—we will continue to assume that all HPs are tuned to maintain notational simplicity.
@@
When choosing a tuning procedure, it is important to consider that the tuning process is limited in terms of both data availability and computation time: First, as outlined above, each candidate HP configuration, $\boldsymbol{\lambda}^{(c)}$, is evaluated using one of the evaluation procedures described in Section 3.2.1 for Setting I. As explained there, the specified $\mathcal{D}^{\prime}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ subsets contain a limited number of observations (i.e. $n^{\prime}_{\text{train}}$ and $n_{\text{test}}$ $\leq n_{\text{train}}$) and could overlap, potentially leading to unreliable prediction error estimates for each $\boldsymbol{\lambda}^{(c)}$. Second, the computational budget available for the tuning process is typically limited, which restricts both the number of evaluated HP configurations and the time spent evaluating each configuration (i.e. estimating its prediction error). Due to these limitations and the resulting trade-offs (discussed in more detail in Section 4.1.3), choosing an adequate tuning procedure is often non-trivial. Yet, guidance is still lacking, and many of the existing recommendations are based on rules of thumb rather than empirical benchmarks (see [78], for an overview). Inadequate tuning procedures can result in a $\boldsymbol{\lambda}^{\mathrm{II}}$ that yields a final prediction model with worse prediction error than $\boldsymbol{\lambda}^{\ast}$ (potentially even worse than setting all HPs to their default values) and/or an overly time-consuming tuning process (i.e. a more efficient tuning procedure could have achieved the same prediction error in less time).
@@
Before describing different tuning procedures in more detail, we note that their specification generally depends on whether the tuning process is fully automated or performed manually. We consider the tuning process as automated if the relevant tuning components only need to be specified as a function argument, which is possible in several ML software frameworks (see [78], for an overview). In contrast, we refer to the tuning process as manual if the candidate HP configurations are evaluated by repeatedly calling the same function(s), altering only the argument that specifies the HP configuration.
Compared to automated tuning, manual tuning is more time-consuming, error-prone, and less reproducible, as it is usually an informal and unsystematic process. On the other hand, automated tuning is usually more difficult to implement and requires more programming expertise than manual tuning. As a consequence, although manual tuning is generally advised against [76, 78, e.g.,], it is likely still a common yet often unreported approach in many ML applications [104, 100, 114]. Note that this may be particularly true for the tuning of preprocessing HPs $\boldsymbol{\lambda}_{P}$: As discussed in Section 2.3.3, preprocessing HPs are often not identified as HPs. Consequently, users trying out different preprocessing options might not be aware that this corresponds to (manual) HP tuning and could be automated. Moreover, if the HPs to be tuned include application-specific preprocessing HPs, the barrier to using automated tuning is further increased, as these HPs may not yet be integrated into the corresponding software and require custom implementation.